# DocKylin: A Large Multimodal Model for Visual Document Understanding with Efficient Visual Slimming

### v3

@@ Introduction
Table 4: Ablation studies on DocKylin regarding APS and DTS. The time of training and inference (Infer.) are normalized to the results in the first row. note that the token length and inference time are obtained based on the DocVQAval dataset. Accuracy proposed in (Zhang et al. 2024) is adopted as the metric for all benchmarks.
APS DTS Token Len. Train Infer. DocVQAval ChartQA
2770 1 1 63.1 55.8
2198 0.97 0.81 64.5 58.5
1250 0.86 0.61 65.1 59.0
To further demonstrate the superiority of our method, we compared it with other parameter-free visual token compression methods, using LLaVA1.5 (Liu et al. 2023a) as our baseline model (without retraining). The results in Table 3 indicate that our APS and DTS achieve performance improvements with shorter sequence lengths compared to the baseline model, significantly outperforming existing SOTA compression methods. It is worth noting that these results are achieved without retraining. Typically, after retraining, models can better adapt to the new mode of compressed token sequences, leading to further performance improvements (Shang et al. 2024; Chen et al. 2024).
Table 5: The Scanned Document, Infographics, Chart, Table, Receipt, and Ingredient List are sourced from the DocVQAval, InfoVQAval, ChartQA (augmented), WTQ,SROIE and POIE datasets, respectively. All benchmarks are evaluated using the accuracy metric as proposed in (Zhang et al. 2024).
@@
Ratio Acc.$\uparrow$ Ratio Acc.$\uparrow$ Ratio Acc.$\uparrow$ Ratio Acc.$\uparrow$ Ratio Acc.$\uparrow$ Ratio Acc.$\uparrow$
@@
DTS-mask 48.2 54.9 35.2 26.3 42.4 73.8 38.1 31.5 44.1 43.2 34.0 25.2
DTS-R-mask 51.8 0.2 64.8 14.6 57.6 4.4 61.9 0.6 55.9 0.1 66.0 0.5
@@
We validate the effectiveness of APS and DTS on DocKylin and present the results in Table 4. Note that, settings without DTS in Table 4 refer to retraining a new model from scratch without employing DTS in the instruction tuning phase, rather than simply removing DTS from the already trained DocKylin model. The results demonstrate that both APS and DTS effectively reduce token sequence length and enhance computational efficiency, while also improving performance. This aligns with findings from previous studies (Liu et al. 2024b), which suggest that excessively long sequences due to high resolution may degrade performance. This degradation likely occurs because the visual sequences contain too many redundant tokens, placing greater demands on the model’s ability to handle long contexts and making the learning more challenging. Furthermore, the APS and DTS methods can also be used during the training phase to accelerate training. Specifically, applying the DTS method during the instruction tuning phase can reduce training time by approximately 15%.
Table 6: The enhancement brought to other MLLMs by Adaptive Pixel Slimming. Relative improvements are highlighted in bold. Accuracy proposed in (Zhang et al. 2024) is adopted as the metric for all benchmarks.
@@
LLaVA1.5 (Liu et al. 2023a) 336$\times$336 8.5 14.7 1.7 0.2
LLaVA1.5 (Liu et al. 2023a) + APS 10.7 (+27.4%) 14.7 (+0%) 3.7 (+118%) 0.92 (+360%)
@@
Monkey (Li et al. 2023b) 896$\times$896 50.1 25.8 41.9 24.1
Monkey (Li et al. 2023b) + APS 56.3 (+12.4%) 27.5 (+6.6%) 47.0 (+12.2%) 27.3 (+13.3%)
We conduct experiments to further explore: 1) How accurately do APS and DTS identify redundant regions? 2) Do APS and DTS demonstrate consistent performance across different data types? Specifically, we apply APS and DTS to identify redundant regions in different document images, and then mask these regions instead of removing them (as shown in Fig. 5). The masked images are used as inputs to an existing MLLM to assess any performance degradation. This approach allows us to isolate the accuracy of APS and DTS from the performance improvements brought by redundancy removal.
@@
Table 5 shows that APS effectively reduces resolution across all image types with minimal performance impact. However, the reduction effects on the table and ingredient list are less pronounced, likely due to the lines spanning the entire image, making APS difficult to identify redundant rows or columns. DTS also effectively reduces resolution across all image types and typically achieves greater resolution reduction than APS, as it can identify local areas beyond entire rows and columns. While DTS may cause noticeable performance drops in the table datasets, our visual analysis suggests this is due to new patterns introduced by masking table lines, not the masking of critical text content. Practically, such potential performance degradation from DTS can be largely mitigated: 1) DTS retains information by similarity-weighted aggregation approach rather than completely discarding redundant tokens; and 2) during image encoding, information from masked areas is integrated into neighboring tokens.
@@
We also conduct additional experiments where we reversed the DTS-mask results by masking the non-redundant regions identified by DTS, denoted as DTS-R-mask. The results of the DTS-R-mask indicate a significant performance drop, further demonstrating the accuracy of DTS in identifying redundant regions.
@@
Since APS is a parameter-free method that requires no additional training or modification of model architecture, it can be easily applied to other MLLMs to enhance their performance on high-resolution document images. To further demonstrate its effectiveness, we apply APS to various existing MLLM models (employing it before the image processing module of these models). As shown in Table 6, APS consistently improves document image tasks across a range of MLLMs with different resolutions.
@@ Conclusions
To address the issues of high input resolution and excessively long token sequences in current MLLMs when counters with VDU tasks, we propose the DocKylin model. It includes an Adaptive Pixel Slimming (APS) preprocessing module that removes redundant regions from document images, effectively reducing their resolution. The model employs a lightweight and flexible image encoding approach to adapt to the varying resolutions and aspect ratios of document image scenarios. Moreover, we propose a token compression module, Dynamic Token Slimming (DTS), which clusters and selects essential tokens from a long visual sequence and employs aggregation to avoid potential information loss. Experiments demonstrate that DocKylin achieves leading performance across multiple VDU benchmarks. Extensive experiments demonstrate that APS and DTS effectively reduce the length of visual tokens, enhancing both computational efficiency and model performance. The parameter-free nature of our APS and DTS modules not only enhances efficiency but also facilitates easy integration into existing multimodal large language models. This feature makes our approach particularly valuable for improving document understanding capabilities in a wide range of applications.
@@ Appendix A Appendix
Table 8: The enhancement brought to other MLLMs by Adaptive Pixel Slimming. Relative improvements are highlighted in bold. Accuracy proposed in (Zhang et al. 2024) is adopted as the metric for all benchmarks.
@@
LLaVA1.5 (Liu et al. 2023a) 336$\times$336 38.7
LLaVA1.5 (Liu et al. 2023a) + APS 40.2 (+4.1%)
@@
Monkey (Li et al. 2023b) 896$\times$896 64.3
Monkey (Li et al. 2023b) + APS 65.2 (+1.4%)
Figure 6: Effects of applying DTS to natural scene images.
@@
Potential Applications for Scene Text Images
@@
The redundant regions in document images are often simple, such as plain backgrounds, distinguishing these regions is relatively easy. This raises the question of whether DTS’s effectiveness is due to this simplicity. Although this paper primarily explores VDU tasks, we are also interested in the generalizability of DTS. Therefore, we conducted some experiments on natural scene images. Specifically, we retrained a scene text image full-text recognition model using some open-source scene text data (Ch’ng and Chan 2017; Singh et al. 2019; Shi et al. 2017; Veit et al. 2016), performing only the first pre-training phase. We then applied DTS to the trained model and tested it on the TextVQA dataset. Some visual results are shown in Fig. 6. It can be seen that DTS is effective in distinguishing nonessential regions even in the complex backgrounds of natural scene images.
@@
We present the visualization results of APS and DTS on different data types in Figure 1. From these visualizations, we can draw the following conclusions: 1) APS effectively removes redundant regions when document backgrounds are simple, but struggles in scenarios involving graphics, document borders, table lines, or photographs. 2) In contrast, DTS is more robust and can handle these challenging cases that APS cannot. 3) While DTS is not perfect and may mistakenly classify some text areas as redundant, the impact of this issue in practical applications is minimal because: 1) As detailed in the methodology section, DTS does not completely discard redundant tokens but retains as much information as possible through similarity-weighted aggregation; 2) Although DTS might mask useful areas at the image level, during the image encoding process, information from these regions is integrated into neighboring tokens.

### v4

@@ Introduction
APS DTS Token Len. Train Infer. DocVQAval ChartQA InfoVQA POIE
2770 1 1 63.1 55.8 34.8 34.5
2198 0.97 0.81 64.5 58.5 34.7 34.9
1250 0.86 0.61 65.1 59.0 34.8 36.1
Table 4: The training and inference time are normalized to the results in the first row. The token length and inference time are obtained based on the DocVQAval dataset. Accuracy from (Zhang et al. 2024b) is adopted as the metric.
To further demonstrate the superiority of our method, we compared it with other parameter-free visual token compression methods, using LLaVA1.5 (Liu et al. 2024b) as our baseline model (without retraining). The results in Table 3 indicate that our APS and DTS achieve performance improvements with shorter sequence lengths than the baseline model, significantly outperforming existing SOTA compression methods. We also found that incorporating weighted aggregation provided a slight performance boost. After model training, we expect enhanced benefits (Shang et al. 2024; Chen et al. 2025), as aggregation introduces an unfamiliar pattern to the model. Theoretically, it offers a higher performance ceiling by retaining more original information.
@@
Mask Ratio Acc.$\uparrow$ Mask Ratio Acc.$\uparrow$ Mask Ratio Acc.$\uparrow$ Mask Ratio Acc.$\uparrow$ Mask Ratio Acc.$\uparrow$ Mask Ratio Acc.$\uparrow$
@@
DTS-mask 48.2 54.9 35.2 26.3 42.4 73.8 38.1 31.5 44.1 43.2 34.0 25.1
Table 5: The Scanned Document, Infographics, Chart, Table, Webpage, Receipt, and Ingredient List are sourced from the DocVQAval, InfoVQAval, ChartQA (augmented), WTQ, visualMRC, SROIE and POIE datasets, respectively. All data, except for the Webpage which uses the CIDEr metric, are evaluated using the accuracy metric as proposed in (Zhang et al. 2024b).
@@
The Effectiveness of APS and DTS on DocKylin. We validate the effectiveness of APS and DTS on DocKylin and present the results in Table 4. Note that, settings without DTS in Table 4 refer to retraining a new model from scratch without employing DTS in the instruction tuning phase, rather than simply removing DTS from the already trained DocKylin model. The results demonstrate that both APS and DTS effectively reduce token sequence length and enhance computational efficiency, while also improving performance in most cases. This aligns with findings from previous studies (Liu et al. 2024d), which suggest that excessively long sequences due to high resolution may degrade performance. This degradation likely occurs because the visual sequences contain too many redundant tokens, placing greater demands on the model’s ability to handle long contexts and making the learning more challenging. Furthermore, the APS and DTS methods can also be used during the training phase to accelerate training. Specifically, applying the DTS method during the instruction tuning phase can reduce training time by approximately 15%.
@@
LLaVA1.5 (Liu et al. 2024b) 336$\times$336 8.5 14.7 1.7 0.2
LLaVA1.5 (Liu et al. 2024b) + APS 10.7 (+27.4%) 14.7 (+0%) 3.7 (+118%) 0.92 (+360%)
@@
Monkey (Li et al. 2024b) 896$\times$896 50.1 25.8 41.9 24.1
Monkey (Li et al. 2024b) + APS 56.3 (+12.4%) 27.5 (+6.6%) 47.0 (+12.2%) 27.3 (+13.3%)
InternVL2 (Chen et al. 2024) 448$\times$448$\times$(1$\sim$12) 76.2 49.5 54.7 41.7
InternVL2 (Chen et al. 2024) + APS 76.1 48.2 54.2 40.6
InternVL2 (Chen et al. 2024) + APS + Resize 77.3 (+1.4%) 49.4 (-0.2%) 55.2 (+0.9%) 43.4 (+4.1%)
Table 6: The enhancement brought to other MLLMs by Adaptive Pixel Slimming. Relative improvements are highlighted in bold. Accuracy proposed in (Zhang et al. 2024b) is adopted as the metric for all benchmarks.
@@
The accuracy of Identified Redundant Regions. We conduct experiments to further explore: 1) How accurately do APS and DTS identify redundant regions? 2) Do APS and DTS demonstrate consistent performance across different data types? Specifically, we apply APS and DTS to identify redundant regions in different document images, and then mask these regions instead of removing them (as shown in Fig. 5). The masked images are used as inputs to an existing MLLM to assess any performance degradation. This approach allows us to isolate the accuracy of APS and DTS from the performance improvements brought by redundancy removal.
@@
Table 5 shows that APS effectively reduces resolution across all image types with minimal performance impact. However, the reduction effects on the table and ingredient list are less pronounced, likely due to the lines spanning the entire image, making APS difficult to identify redundant rows or columns. DTS also effectively reduces resolution across all image types and typically achieves greater resolution reduction than APS, as it can identify local areas beyond entire rows and columns. While DTS may cause noticeable performance drops in the table datasets, our visual analysis suggests this is due to new patterns introduced by masking table lines, not the masking of critical text content. Practically, such potential performance degradation from DTS can be largely mitigated: 1) DTS retains information by similarity-weighted aggregation approach rather than completely discarding redundant tokens; and 2) during image encoding, information from masked areas is already integrated into neighboring tokens.
@@
We conduct additional experiments by reversing the APS-mask and DTS-mask results by masking the non-redundant regions identified by APS and DTS, denoted as APS-R-mask and DTS-R-mask. The results indicate a significant performance drop, further demonstrating the accuracy of APS and DTS in identifying redundant regions.
We further compare our approach with methods that directly select redundant tokens based on similarity priors (e.g., TextMonkey and PruMerge adopt this approach). Specifically, in Table 5, Top-30-mask and Top-40-mask represent masking the top 30% and 40% of tokens with the highest similarity to other tokens, respectively, as nonessential tokens. Results show that DTS outperforms the top K strategy, even at higher reduction rates.
@@
Potentials for Further Application and Research. Since APS is a parameter-free method that requires no additional training or modification of model architecture, it can be easily applied to other MLLMs to enhance their performance on high-resolution document images. To further demonstrate its effectiveness, we apply APS to various existing MLLMs (employing it before the image processing module of these models). As shown in Table 6, APS consistently improves the performance of MLLMs that support lower resolutions (LLaVA1.5, Qwen-VL and Monkey). However, directly applying APS to high-resolution models like InternVL2 results in a slight performance decline, likely due to a mismatch between the dense text produced by APS and the high-resolution training setup of InternVL2. Adjusting the APS-processed images back to their original resolution (InternVL2+APS+Resize) can achieve a performance gain of up to 4.1%. This suggests that APS alone is most effective for low-resolution models, while APS+Resize is more suitable for high-resolution models.
Figure 6: Effects of applying DTS to natural scene images.
The redundant regions in document images are often simple, such as plain backgrounds, distinguishing these regions is relatively easy. This raises the question of whether DTS’s effectiveness is due to this simplicity. Although this paper primarily explores VDU tasks, we are also interested in the generalizability of DTS. Therefore, we conducted some experiments on scene text images (Tang et al. 2022; Liu et al. 2023b; Zhang et al. 2020). Specifically, we retrained a scene text image full-text recognition model using some open-source scene text data, performing only the first pre-training phase. We then applied DTS to the trained model and tested it on the TextVQA dataset. Some visual results are shown in Fig. 6. It can be seen that DTS is effective in distinguishing nonessential regions even in the complex backgrounds of natural scene images.
@@ Conclusions
To address the issues of high input resolution and excessively long token sequences in current MLLMs when counters with VDU tasks, we propose the DocKylin model. It includes an Adaptive Pixel Slimming (APS) preprocessing module that removes redundant regions from document images, effectively reducing their resolution. The model employs a lightweight and flexible image encoding approach to adapt to the varying resolutions and aspect ratios of document image scenarios. Moreover, we propose a token compression module, Dynamic Token Slimming (DTS), which clusters and selects essential tokens from a long visual sequence and employs aggregation to avoid potential information loss. Experiments demonstrate that DocKylin achieves leading performance across multiple VDU benchmarks. Extensive experiments demonstrate that APS and DTS effectively reduce the length of visual tokens, enhancing both computational efficiency and model performance. The parameter-free nature of our APS and DTS modules not only enhances efficiency but also facilitates easy integration into existing multimodal large language models, and our experiments indicate their potential for broader applications and scenarios. This feature makes our approach particularly valuable for improving document understanding capabilities in a wide range of applications.
Acknowledgments
This research is supported in part by the National Natural Science Foundation of China (Grant No.: 62441604, 62476093).
@@ Appendix A Appendix
Table 8: The enhancement brought to other MLLMs by Adaptive Pixel Slimming. Relative improvements are highlighted in bold. Accuracy proposed in (Zhang et al. 2024b) is adopted as the metric for all benchmarks.
@@
LLaVA1.5 (Liu et al. 2024b) 336$\times$336 38.7
LLaVA1.5 (Liu et al. 2024b) + APS 40.2 (+4.1%)
@@
Monkey (Li et al. 2024b) 896$\times$896 64.3
Monkey (Li et al. 2024b) + APS 65.2 (+1.4%)
@@
APS’s Performance for Scene Text Images
@@
We present the visualization results of APS and DTS on different data types in Figure 1. From these visualizations, we can draw the following conclusions: 1) APS effectively removes redundant regions when document backgrounds are simple, but struggles in scenarios involving graphics, document borders, table lines, or photographs. 2) In contrast, DTS is more robust and can handle these challenging cases that APS cannot. 3) While DTS is not perfect and may mistakenly classify some text areas as redundant, the impact of this issue in practical applications is minimal because: As detailed in the methodology section, DTS does not completely discard redundant tokens but retains as much information as possible through similarity-weighted aggregation; and although DTS might mask useful areas at the image level, during the image encoding process, information from these regions is integrated into neighboring tokens.
@@
Edge Cases for DTS
As shown in Algorithm 1, it is possible for DTS to encounter a situation where num1 equals num2. To investigate this issue, we re-evaluated the DocVQA dataset, recording values for num1 and num2 across 4335 samples. We found that the maximum, minimum, mean, and variance of num1 - num2 are 50, 0, 47, and 9, respectively (with num1 + num2 = 50). Only one sample had num1 - num2 = 0, and the proportion of samples with a difference of less than 10 was only 0.2%. These indicate that num1 - num2 = 0 is a very rare case, with high confidence in most instances (large differences). Further analysis of this rare sample revealed a low input resolution with minimal redundant areas, leading some non-redundant tokens to be mistakenly included in the top 50 tokens by similarity. This issue can be resolved by adding a check to bypass DTS when the confidence is low (small difference).